# Dingjie Song, Shunian Chen, Guiming Hardy Chen, Fei Yu, Xiang Wan, Benyou Wang

### v1

@@ Abstract
Despite the advancements and impressive performance of Multimodal Large Language Models (MLLMs) on benchmarks, their effectiveness in real-world, long-context, and multi-image tasks is unclear due to the benchmarks’ limited scope. Existing benchmarks often focus on single-image and short-text samples, and when assessing multi-image tasks, they either limit the image count or focus on specific task (e.g time-series captioning), potentially obscuring the performance challenges of MLLMs. To address these limitations, we introduce MileBench, a pioneering benchmark designed to test the MultImodal Long-contExt capabilities of MLLMs. This benchmark comprises not only multimodal long contexts, but also multiple tasks requiring both comprehension and generation. We establish two distinct evaluation sets, diagnostic and realistic, to systematically assess MLLMs’ long-context adaptation capacity and their ability to complete tasks in long-context scenarios. Our experimental results, obtained from testing 20 models, revealed that while the closed-source GPT-4(Vision) and Gemini 1.5 outperform others, most open-source MLLMs struggle in long-context situations. Interestingly, the performance gap tends to widen with an increase in the number of images. We strongly encourage an intensification of research efforts towards enhancing MLLMs’ long-context capabilities, especially in scenarios involving multiple images.
@@ 1 Introduction
After evaluating 20 models, the closed-source Gemini 1.5 (Reid et al., 2024) excelled in the realistic evaluation, achieving an impressive score of 54.7%, though it still falls short of a perfect 100% score. Meanwhile, GPT-4(Vision) (OpenAI, 2023) managed to reach a peak score of 99.4% in the diagnostic evaluation. On the contrary, most open-source MLLMs struggled with long-context tasks as depicted in Figure 1. Only VILA and Qwen-VL-7B managed average scores of 44.4% and 37.2% in realistic and diagnostic evaluations respectively. These results underscore that there are “miles to go” towards fully-realized long-context MLLMs, prompting a call for increased research focus on such tasks, especially those involving numerous images.
@@ 4.1 Experiment Setup
In this study, we conducted an evaluation of several models across three distinct categories that may handle multimodal long contexts, including four closed-source models [^5] Since the performance of closed-source models may change over time, we report testing GPT-4V (gpt-4-turbo-2024-04-09) on April 12, 2024, Claude 3 (claude-3-opus-20240229) on March 16, 2024, Gemini 1.0 (gemini-1.0-pro-vision-001) on March 20, 2024, and Gemini 1.5 (gemini-1.5-pro-latest) on April 12, 2024. Please note that Gemini 1.0 has a limitation of accepting up to 16 images as input and Claude 3 Opus is 20. (GPT-4V (OpenAI, 2023), Gemini 1.0 (Anil et al., 2023), Gemini 1.5 (Reid et al., 2024), Claude 3 Opus [^6] https://www.anthropic.com/news/claude-3-family), eleven open-source image models (Qwen-VL-Chat (Bai et al., 2023), MiniGPT-v2 (Chen et al., 2023), Cheetor (Li et al., 2023c), Open flamingo (Awadalla et al., 2023), LLaVA-1.5-7B/13B (Liu et al., 2023a), LLaVA-1.6-7B/13B (Liu et al., 2024b), ALLaVA-Longer (Chen et al., 2024), Yi-VL (Young et al., 2024)), VILA (Lin et al., 2023)), and five open-source video models (Video-LLaMA-2 (Zhang et al., 2023), Valley (Luo et al., 2023), VideoChat2 (Li et al., 2023d), LLaMA-VID (Li et al., 2023e), LWM (Liu et al., 2024a)). The details of the models are in Appendix C.1. All models used greedy decoding to generate answers, with a designated generation length between 1 and 512. We conducted all experiments on NVIDIA A100 GPUs.
@@ 4.2.1 Main Result on MileBench
We present the results of our experiments in Table 2 and summarize our findings as follows: (1) Closed-source MLLMs outperform open-source MLLMs in multimodal long-context tasks to date, particularly in diagnostic evaluation of long-context adaptability where the gap between closed-source MLLMs (average: 74.2%, max: 99.4%) and open-source MLLMs (average: 9.5%, max: 37.2%) is significant. In realistic evaluation, all open-source models, except for VILA, lag considerably behind. (2) Open-source image models generally perform better than open-source video models. Even the best video model, LLaMA-VID, falls short in realistic evaluation with 31.8%, a score that is lower than seven image models. This may be due to the inability of video models to capture detailed information in images in the same way that image models can. (3) The ability to adapt to long contexts and perform long-context tasks are not necessarily linked. For example, while Qwen-VL-Chat scores the highest in diagnostic evaluation among open-source models, it trails behind VILA in task completion (39.1% <44.4%), highlighting our evaluation’s diversity and comprehensiveness. (4) Interestingly, the majority of open-source models scored zero in the Image Needle in a Haystack task. Upon inspection, we found that many of these models partially answered the needle numeric string without completely getting it right. This suggests that open-source models need to improve their ability to retrieve information from images, particularly their OCR capabilities. Detailed results from the realistic evaluation can be found in Appendix C.3.
@@ 4.3 Analysis
    Do MLLMs get “Lost in the Middle” (Liu et al., 2023b) in multimodal long contexts?
@@
    Does combining image help multi-image comprehension?
@@ 4.3.1 Performances Decline as the Number of Images Increases for Most MLLMs
To investigate the performance of MLLMs with varying numbers of images, we divide our dataset into three levels: Few, Medium, and Many, based on the number of images per sample. The specific quantities for each level can be found in Table 4. Figure 6 reports the average performance of the model on the three types of data with different numbers of images. It can be observed that as the number of images increases, the performance of most models significantly declines (as indicated by a steep slope in the curve), especially for the LLaVA-1.5 series models. This is likely because most models have only been trained on single image, resulting in insufficient generalization for multi-image test data. However, the performance of GPT-4V, Gemini 1.5, Claude 3 Opus and Qwen-VL-Chat on the Medium level surpasses that of the Few level. This could be attributed to their training on multi-image data, where a larger number of images can provide more information to some extent, thereby aiding the model in task completion. Despite their outstanding performance on multi-image tasks, their performance still declines when the number of images reaches the Many level. This leaves room for future development in modeling for multi-image context.
@@
Some literature pointed out that in needle-in-a-haystack tasks involving long texts, models may experience the “Lost in the Middle” phenomenon, where they struggle to find the needle located in the middle of the context. We investigated whether MLLMs would exhibit the “Lost in the Middle” phenomenon in multimodal contexts. We chose the two best-performing models from closed-source and open-source models in the Needle in a Haystack task for analysis. As can be seen from the results in Figure 7, MLLMs displayed varying behaviors. In multimodal long contexts, GPT-4V did not “get lost in the middle” and managed to complete the two tasks impressively with the scores 99.7% (N-1) and 99.1% (N-2). On the other hand, ignoring the scenarios where the data exceeds its maximum context length (8192 tokens or 32 images) and gets truncated, Qwen-VL-Chat showed a certain degree of “lost in the middle”, particularly evident in the image needle task. This indicates that the “lost in the middle” phenomenon also exists in multimodal scenarios. However, a strong ability to manage long context can significantly reduce this risk.
@@ 4.3.4 Experiment on Combined-image Set
We show the result on combined image set in Table 4 and summarize our findings as follows: (1) The performance of proprietary models still surpasses that of open-source models in both realistic evaluation (average: 44.8% v.s. 28.6%, max: 48% v.s. 41.9%) and diagnostic evaluation (average: 51.2% v.s. 11.0%, max: 60.5% v.s. 20.6%). (2) In comparison to the results on the multi-image set, the performance of proprietary models declined, except for Gemini 1.0, which is limited by the number of images uploaded on the multi-image set. The potential reason is that to maintain performance on the combined image set, models need to possess high-resolution vision models that can effectively distinguish multiple images combined together. For instance, Gemini 1.0 adjusts images of excessively large resolution to a size of $3072\times 3072$. On the other hand, GPT-4V and Claude 3 Opus resize the images to dimensions of $768\times 768$ and $1568\times 1568$, respectively. The lower resolution input of GPT-4V and Claude 3 Opus, in comparison to Gemini 1.0, could potentially be a factor for their diminished performance. (3) Compared to the results on the multi-image set, the performance of some open-source models with short contexts improved, such as ALLaVA-Longer (from 24.7% to 26.9%) and MiniGPT-v2 (from 17.8% to 29.5%). The possible reason is that these models have only been trained on single images and cannot effectively generalize to multi-image scenarios. Combining images can effectively alleviate this issue.
@@ C.1 Details of Evaluation Models
  • GPT-4V (OpenAI, 2023) is developed by OpenAI and is deemed the most powerful vision-language model for comprehension and generation.
  • Gemini 1.0 Pro (Anil et al., 2023) is a model developed by Google, demonstrating competitive ability among closed-source MLLMs.

### v2

@@ Abstract
Despite the advancements and impressive performance of Multimodal Large Language Models (MLLMs) on benchmarks, their effectiveness in real-world, long-context, and multi-image tasks is unclear due to the benchmarks’ limited scope. Existing benchmarks often focus on single-image and short-text samples, and when assessing multi-image tasks, they either limit the image count or focus on specific task (e.g time-series captioning), potentially obscuring the performance challenges of MLLMs. To address these limitations, we introduce MileBench, a pioneering benchmark designed to test the MultImodal Long-contExt capabilities of MLLMs. This benchmark comprises not only multimodal long contexts, but also multiple tasks requiring both comprehension and generation. We establish two distinct evaluation sets, diagnostic and realistic, to systematically assess MLLMs’ long-context adaptation capacity and their ability to complete tasks in long-context scenarios. Our experimental results, obtained from testing 22 models, revealed that while the closed-source GPT-4o outperforms others, most open-source MLLMs struggle in long-context situations. Interestingly, the performance gap tends to widen with an increase in the number of images. We strongly encourage an intensification of research efforts towards enhancing MLLMs’ long-context capabilities, especially in scenarios involving multiple images.
@@ 1 Introduction
After evaluating 22 models, the closed-source GPT-4o [^2] https://openai.com/index/hello-gpt-4o excelled in both diagnostic and realistic evaluations, achieving impressive scores of 99.4% and 60.3%, although it still falls short of a perfect 100% score. On the contrary, most open-source MLLMs struggled with long-context tasks as depicted in Figure 1. Only Mantis and Qwen-VL-7B managed average scores of 47.5% and 37.2% in realistic and diagnostic evaluations respectively. These results underscore that there are “miles to go” towards fully-realized long-context MLLMs, prompting a call for increased research focus on such tasks, especially those involving numerous images.
@@ 4.1 Experiment Setup
In this study, we conducted an evaluation of several models across three distinct categories that may handle multimodal long contexts, including five closed-source models [^6] Since the performance of closed-source models may change over time, we report testing GPT-4V (gpt-4-turbo-2024-04-09) on April 12, 2024, GPT-4o (gpt-4o-2024-05-13) on May 13, 2024, Claude 3 (claude-3-opus-20240229) on March 16, 2024, Gemini 1.0 (gemini-1.0-pro-vision-001) on March 20, 2024, and Gemini 1.5 (gemini-1.5-pro-latest) on April 12, 2024. Please note that Gemini 1.0 has a limitation of accepting up to 16 images as input and Claude 3 Opus is 20. (GPT-4V (OpenAI, 2023), GPT-4o, Gemini 1.0 (Anil et al., 2023), Gemini 1.5 (Reid et al., 2024), Claude 3 Opus [^7] https://www.anthropic.com/news/claude-3-family), twelve open-source image models (Qwen-VL-Chat (Bai et al., 2023), MiniGPT-v2 (Chen et al., 2023), Cheetor (Li et al., 2023c), Open flamingo (Awadalla et al., 2023), LLaVA-1.5-7B/13B (Liu et al., 2023a), LLaVA-1.6-7B/13B (Liu et al., 2024b), ALLaVA-Longer (Chen et al., 2024), Yi-VL (Young et al., 2024), VILA (Lin et al., 2023), Mantis (Jiang et al., 2024)), and five open-source video models (Video-LLaMA-2 (Zhang et al., 2023), Valley (Luo et al., 2023), VideoChat2 (Li et al., 2023d), LLaMA-VID (Li et al., 2023e), LWM (Liu et al., 2024a)). The details of the models are in Appendix C.1. All models used greedy decoding to generate answers, with a designated generation length between 1 and 512. We conducted all experiments on NVIDIA A100 GPUs.
@@ 4.2.1 Main Result on MileBench
We present the results of our experiments in Table 2 and summarize our findings as follows: (1) Closed-source MLLMs outperform open-source MLLMs in multimodal long-context tasks to date, particularly in diagnostic evaluation of long-context adaptability where the gap between closed-source MLLMs (average: 79.2%, max: 99.4%) and open-source MLLMs (average: 10.1%, max: 37.2%) is significant. In realistic evaluation, all open-source models, except for VILA and Mantis, lag considerably behind. (2) Open-source image models generally perform better than open-source video models. Even the best video model, LLaMA-VID, falls short in realistic evaluation with 31.8%, a score that is lower than eight image models. This may be due to the inability of video models to capture detailed information in images in the same way that image models can. (3) The ability to adapt to long contexts and perform long-context tasks are not necessarily linked. For example, while Qwen-VL-Chat scores the highest in diagnostic evaluation among open-source models, it trails behind Mantis in task completion (39.1% <47.5%), highlighting our evaluation’s diversity and comprehensiveness. (4) Interestingly, the majority of open-source models scored zero in the Image Needle in a Haystack task. Upon inspection, we found that many of these models partially answered the needle numeric string without completely getting it right. This suggests that open-source models need to improve their ability to retrieve information from images, particularly their OCR capabilities. Detailed results from the realistic evaluation can be found in Appendix C.3.
@@ 4.3 Analysis
    Do MLLMs also get “Lost in the Middle” in multimodal long contexts?
@@
    Does combined image help multi-image comprehension?
@@ 4.3.1 Performances Decline as the Number of Images Increases for Most MLLMs
To investigate the performance of MLLMs with varying numbers of images, we divide our dataset into three levels: Few, Medium, and Many, based on the number of images per sample. The specific quantities for each level can be found in Table 4. Figure 6 reports the average performance of the model on the three types of data with different numbers of images. It can be observed that as the number of images increases, the performance of most models significantly declines (as indicated by a steep slope in the curve), especially for the LLaVA-1.5 series models. This is likely because most models have only been trained on single image, resulting in insufficient generalization for multi-image test data. However, the performance of GPT-4V, GPT-4o, Gemini 1.5, Claude 3 Opus and Qwen-VL-Chat on the Medium level surpasses that of the Few level. This could be attributed to their training on multi-image data, where a larger number of images can provide more information to some extent, thereby aiding the model in task completion. Despite their outstanding performance on multi-image tasks, their performance still declines when the number of images reaches the Many level. This leaves room for future development in modeling for multi-image context.
@@
Liu et al. (2023b) pointed out that in needle-in-a-haystack tasks involving long texts, models may experience the “Lost in the Middle” phenomenon, where they struggle to find the needle located in the middle of the context. We investigated whether MLLMs would exhibit the “Lost in the Middle” phenomenon in multimodal contexts. We chose the two best-performing models from closed-source and open-source models in the Needle in a Haystack task for analysis. As can be seen from the results in Figure 7, MLLMs displayed varying behaviors. In multimodal long contexts, GPT-4V did not “get lost in the middle” and managed to complete the two tasks impressively with the scores 99.7% (N-1) and 99.1% (N-2). On the other hand, ignoring the scenarios where the data exceeds its maximum context length (8192 tokens or 32 images) and gets truncated, Qwen-VL-Chat showed a certain degree of “lost in the middle”, particularly evident in the image needle task. This indicates that the “lost in the middle” phenomenon also exists in multimodal scenarios. However, a strong ability to manage long context can significantly reduce this risk.
@@ 4.3.4 Experiment on Combined-image Set
We show the result on combined image set in Table 4 and summarize our findings as follows: (1) The performance of proprietary models still surpasses that of open-source models in both realistic evaluation (average: 44.8% v.s. 29.6%, max: 48% v.s. 44.9%) and diagnostic evaluation (average: 51.2% v.s. 12.3%, max: 60.5% v.s. 32.4%). (2) In comparison to the results on the multi-image set, the performance of proprietary models declined, except for Gemini 1.0, which is limited by the number of images uploaded on the multi-image set. The potential reason is that to maintain performance on the combined image set, models need to possess high-resolution vision models that can effectively distinguish multiple images combined together. For instance, Gemini 1.0 adjusts images of excessively large resolution to a size of $3072\times 3072$. On the other hand, GPT-4V and Claude 3 Opus resize the images to dimensions of $768\times 768$ and $1568\times 1568$, respectively. The lower resolution input of GPT-4V and Claude 3 Opus, in comparison to Gemini 1.0, could potentially be a factor for their diminished performance. (3) Compared to the results on the multi-image set, the performance of some open-source models with short contexts improved, such as ALLaVA-Longer (from 24.7% to 26.9%) and MiniGPT-v2 (from 17.8% to 29.5%). The possible reason is that these models have only been trained on single images and cannot effectively generalize to multi-image scenarios. Combining images can effectively alleviate this issue.
@@ C.1 Details of Evaluation Models
  • GPT-4V (OpenAI, 2023) and GPT-4o [^8] https://openai.com/index/hello-gpt-4o/ are developed by OpenAI and are deemed the most powerful vision-language models for comprehension and generation.
  • Gemini 1.0 (Anil et al., 2023) and Gemini 1.5 (Reid et al., 2024) are models developed by Google, demonstrating competitive ability among closed-source MLLMs.
@@
  • Mantis (Jiang et al., 2024) uses LLaMA3101010https://github.com/meta-llama/llama3 as language model and is trained with multi-image data.